# Can Video Large Multimodal Models Think Like Doubters—or Double-Down: A Study on Defeasible Video Entailment

### v1

@@ Appendix C Dataset Comparison
Existing video understanding benchmarks like Charades [Sigurdsson et al., 2016], InfiniBench [Ataallah et al., 2024], and Kinetics [Kay et al., 2017] provide extensive evaluations of model abilities to interpret and reason about fixed, predefined video scenes, such as recognizing actions, classifying events, and understanding temporal relationships. However, they do not assess a model’s ability to handle dynamic and nuanced semantic changes introduced by new, uncertain information In contrast, DVidE explicitly evaluates a model’s ability to adapt its reasoning when presented with new evidence.
Natural Language Inference (NLI) datasets assess whether a hypothesis follows from a premise, typically classifying their relationship as entailment, contradiction, or neutral. Datasets such as SNLI [Bowman et al., 2015] and MNLI [Williams et al., 2018] focus solely on textual entailment without considering inference updates. To introduce defeasibility, datasets like $\delta$-NLI [Rudinger et al., 2020] and $\delta$-CAUSAL [Cui et al., 2024] provide updates that strengthen or weaken inference. However, these defeasible NLI datasets remain text-only and do not handle multi-modal reasoning. Furthermore, DVE [Zhang et al., 2024] introduces image-based defeasible reasoning by replacing the textual premise with an image. However, it remains limited to single images and lacks temporal reasoning. In contrast, DVidE explicitly introduces defeasibility in videos, allowing models to revise inferences as new temporal visual and auditory information flows in.
Appendix D Case Study
@@
Appendix E Limitations

### v2

@@ Appendix C Dataset Comparison
Existing video understanding benchmarks like Charades (Sigurdsson et al., 2016), InfiniBench (Ataallah et al., 2024), and Kinetics (Kay et al., 2017) provide extensive evaluations of model abilities to interpret and reason about fixed, predefined video scenes, such as recognizing actions, classifying events, and understanding temporal relationships. However, they do not assess a model’s ability to handle dynamic and nuanced semantic changes introduced by new, uncertain information In contrast, DVidE explicitly evaluates a model’s ability to adapt its reasoning when presented with new evidence.
Natural Language Inference (NLI) datasets assess whether a hypothesis follows from a premise, typically classifying their relationship as entailment, contradiction, or neutral. Datasets such as SNLI (Bowman et al., 2015) and MNLI (Williams et al., 2018) focus solely on textual entailment without considering inference updates. To introduce defeasibility, datasets like $\delta$-NLI (Rudinger et al., 2020) and $\delta$-CAUSAL (Cui et al., 2024) provide updates that strengthen or weaken inference. However, these defeasible NLI datasets remain text-only and do not handle multi-modal reasoning. Furthermore, DVE (Zhang et al., 2025) introduces image-based defeasible reasoning by replacing the textual premise with an image. However, it remains limited to single images and lacks temporal reasoning. In contrast, DVidE explicitly introduces defeasibility in videos, allowing models to revise inferences as new temporal visual and auditory information flows in.
Appendix D Ethics Statement
D.1 Human Annotation Protocol
Recruitment & Screening. We employed paid crowdworkers and graduate research assistants via an internal university pool. Candidates completed a qualification task with 20 multiple–choice items based on exemplar annotations; only those scoring $\geq 85\%$ advanced to production.
Compensation. Annotators were paid \$20/hour, above the local minimum wage and aligned with institutional guidance for human–subjects research compensation.
Instructions & Protections. Annotators received written guidelines describing the task (identify or compose strengthener/weakener updates for a given video–hypothesis pair), edge cases, and quality criteria. No personally identifying information (PII) was collected.
D.2 Data Provenance, Licensing, and Intended Use
Our benchmark augments VIOLIN video–hypothesis pairs with defeasible updates. We do not collect new private media; instead, we reference timestamped clips retrievable from the original VIOLIN sources for transparency and traceability. Released materials (textual updates/labels and retrieval metadata) are for non–commercial research only; users must comply with the terms and licenses of the underlying sources.
D.3 Privacy, Safety, and Content Risk
The dataset contains no PII (faces/voices occur within publicly released TV/movie material in VIOLIN). Because scripted media may include harmful or offensive themes, we provide a content disclaimer and advise against use in safety–critical or sensitive settings (e.g., hiring, policing, healthcare). Downstream users should conduct separate safety and bias assessments before any deployment.
D.4 Release Protocol
We will release our data after this work has been accepted, under a Creative Commons Attribution–NonCommercial 4.0 International (CC BY-NC 4.0) license. This license permits redistribution and adaptation for non-commercial research purposes with appropriate credit.
Appendix E Case Study
@@
Appendix F Usage of LLMs
We made limited use of large language models—specifically ChatGPT and GitHub Copilot—for language polishing and routine coding assistance (autocompletion, refactoring, and boilerplate generation). LLMs were not used to generate research ideas, experimental results, or claims. No non-public or sensitive data was shared with these tools. All suggested text and code were independently reviewed, tested, and edited by the authors, who remain responsible for the content.